If the perimeter of a rectangle is 104 inches and the length of the rectangle is 6 inches less than its width, find the area of the rectangle.
Answer by mananth(16949)

width = w
length = w-6
..
perimeter = 2(w+l)
Perimeter =2(w+w-6)
Perimeter =4w-12
4w-12=104
+12
4w=116
/4
w=29 inches
length = 29-6=23 inches
...